# Learning Network Representations with Disentangled Graph Auto-Encoder

### v1

@@ Abstract
The (variational) graph auto-encoder is extensively employed for learning representations of graph-structured data. However, the formation of real-world graphs is a complex and heterogeneous process influenced by latent factors. Existing encoders are fundamentally holistic, neglecting the entanglement of latent factors. This not only makes graph analysis tasks less effective but also makes it harder to understand and explain the representations. Learning disentangled graph representations with (variational) graph auto-encoder poses significant challenges, and remains largely unexplored in the existing literature. In this article, we introduce the Disentangled Graph Auto-Encoder (DGA) and Disentangled Variational Graph Auto-Encoder (DVGA), approaches that leverage generative models to learn disentangled representations. Specifically, we first design a disentangled graph convolutional network with multi-channel message-passing layers, as the encoder aggregating information related to each disentangled latent factor. Subsequently, a component-wise flow is applied to each channel to enhance the expressive capabilities of disentangled variational graph auto-encoder. Additionally, we design a factor-wise decoder, considering the characteristics of disentangled representations. In order to further enhance the independence among representations, we introduce independence constraints on mapping channels for different latent factors. Empirical experiments on both synthetic and real-world datasets show the superiority of our proposed method compared to several state-of-the-art baselines.
@@
Keywords: Graph auto-encoder, disentangled representation learning
1 Introduction
@@
Graph-structured data has become exceptionally prevalent in the era of big data, spanning across various domains, including social networks, co-authorship networks, biological networks and traffic networks, etc. Represented in the form of graphs, these data illustrate the complex interconnections among entities, reflecting intricate and extensive networks of relationships.
The field of graph analysis includes many significant tasks. Link prediction (Loehlin, 2004; Wang et al., 2008; Daud et al., 2020; Wang et al., 2021; Son et al., 2023) is one of the best-known tasks for predicting missing or possible connections within a graph. For example, in social networks, link prediction helps with precise targeting in target marketing, which can guide marketing strategies. In biological networks, it is utilized to predict connections between molecules, while in terrorist networks, it assists in identifying suspicious individuals. Other popular graph analysis tasks include node clustering (Wang et al., 2021; Tsitsulin et al., 2023) and node classification (Bhagat et al., 2011; Xiao et al., 2022; Luan et al., 2023). Node clustering aims to group nodes in a graph, where nodes in each cluster are highly relevant. This task finds extensive applications across various domains, including social networks, biology, network security, logistics, finance, and medical image analysis. For instance, in biological networks, it can identify functional modules or groups of genes that work together. Node classification involves assigning predefined labels to nodes based on their attributes, and this task is crucial for understanding the characteristics of individual nodes within a graph. For instance, it can help predict the topics or fields of research papers in citation networks.
However, due to the extensive and highly sparse nature of graphs, graph analysis tasks are often challenging. Therefore, there has been a growing interest in enhancing the learning of graph representations (Perozzi et al., 2014; Hamilton et al., 2017; Wang et al., 2016, 2017; Goyal et al., 2018; Wang et al., 2022). The node representation learning tasks form the foundation of the network analysis tasks mentioned earlier. One particularly successful model for unsupervised node representation learning is the Variational Graph Auto-Encoder (VGAE) and the Graph Auto-Encoder (GAE) (Kipf et al., 2016a). They all consist of an encoder responsible for learning latent representations and a decoder tasked with reconstructing the graph structure data using the learned representations. Their primary goal is to accomplish the task of link prediction. As a result, the decoder is relatively straightforward, relying solely on the inner product of the latent representations of the nodes involved for link predictions. However, the encoder is only built upon the original Graph Convolutional Networks (GCN) (Kipf et al., 2016b).
Figure 1: An illustrative social network example that inspired our work. We consider the reasons for interactions between two individuals annotated with lines of different colors, where the thickness of the lines represents the primary determining factor for interaction frequency.
Figure 2: The whole framework of the proposed DVGA. It takes in nodes and their neighbors along with their feature vectors. The dynamic disentangled encoder utilizes DGCN, where the dynamic assignment mechanism considers neighborhoods induced by $K$ different factors and additionally accounts for neighborhoods in the $k$-th component space. After channel-wise convolution, the resulting component representations are merged and fed into flows for learning disentangled node representations, which are then input into a factor-wise decoder to reconstruct the adjacency matrix. A classifier $g_{\phi}$ is employed to promote the independence of disentangled representations across different latent factors. The joint optimization of the factor-wise auto-encoder and independence regularization enhances the disentanglement. In this example, we assume the existence of three latent factors, each corresponding to one of the three channels.
Despite the success of auto-encoder models, existing methods often employ GCN in the encoder, taking a holistic approach and treating the graph as a perceptual whole (Kipf et al., 2016a; Hasanzadeh et al., 2019; Grover et al., 2019; Khan et al., 2021; Shi et al., 2020; Sun et al., 2021; Li et al., 2023). They ignore subtle differences between various relationships in the graph. In fact, graph formation in the real world is a complex, heterogeneous process driven by latent factors. Figure 1 illustrates the concept of disentanglement. It is evident that this social network is inherently heterogeneous, with connections between the four individuals formed for different underlying factors. Moreover, their interactions exhibit distinct emphases. For instance, the interaction between person $A$ and person $B$ is primarily due to familial ties, while person $A$ and person $C$ engage more due to a shared hobby. Additionally, it’s worth noting that connections between every pair of individuals occur due to multiple factors, but the emphasis varies. Without this fine-grained distinction in relationships, representing nodes only from binary graph structures becomes challenging. This not only hinders the interpretability of learned representations, but also lead to reduced performance of downstream tasks. Therefore, this complexity urges the encoder to disentangle factors, an aspect overlooked by the existing holistic approaches.
@@
In this paper, we propose learning disentangled node representations. While disentangled representation learning, which aims to characterize various underlying factors in factorized representations, has shown many advantages (Higgins et al., 2018), it faces three challenges in graph auto-encoder models. (1) Designing a graph encoder for disentangled representations requires careful planning to guarantee it can effectively infer latent factors with enough expressiveness. (2) Designing a decoder for disentangled representations involves considering factor information in node connections to improve link prediction accuracy. (3) Designing a training scheme to enhance representation independence involves establishing statistical independence between different parts of representations corresponding to different latent factors, thereby improving the quality of disentangled representations.
@@
To tackle the aforementioned challenges, we introduce innovative auto-encoder models for learning disentangled graph representations, that is, Disentangled Variational Graph Auto-Encoder (DVGA) and Disentangled Auto-Encoder (DGA). Initially, we design a Disentangled Graph Convolutional Network (DGCN) using dynamic assigning mechanism (Zheng et al., 2021), incorporating multiple channels to aggregate information related to each disentangled latent factor. It facilitates the learning of disentangled latent representations. Subsequently, a component-wise flow is applied to each channel to enhance the expressiveness of individual representations, promoting the disentangled representations to best capture representations related to latent factors in the graph. Furthermore, we design a factor-wise decoder, ensuring that if there is any factor leading to a connection between nodes, then these two nodes are connected in the final prediction result. This decoder improves the performance of downstream tasks. Additionally, to encourage the independence of representations corresponding to different latent factors, we introduce independence constraints to the mapping channels of different latent factors. Finally, the auto-encoder and independent regularization will be co-optimized within a unified framework, empowering the disentangled graph encoder to generate superior disentangled graph representations while enhancing overall model performance. In comparison to existing methods, our models learn disentangled representations from the graph, allowing for exploration of the meaning of each channel, thereby improving interpretability and generating convincing and understandable link prediction results.
@@
  • $\bullet$
    We introduce two novel disentangled graph auto-encoder models: the Disentangled Variational Graph Auto-Encoder (DVGA) and the Disentangled Graph Auto-Encoder (DGA). The DVGA model incorporates a disentangled encoder, component-wise flow, factor-wise decoder, and an independence regularization term in the loss function to learn disentangled representations. Simultaneously, the DGA model is similarly presented. To the best of our knowledge, we are among the very few researchers investigating disentangled graph auto-encoder models.
  • $\bullet$
    Our proposed disentangled graph convolutional network in the encoder captures multiple factors by learning disentangled latent factors on graphs, enabling multi-relation disentangling. We enhance the expressive capability of each channel representation through flow, further boosting the overall capabilities of the representation. The designed decoder is tailored for disentangled representations, improving its ability to predict graph structures.
  • $\bullet$
@@
The subsequent sections of the paper are structured as follows. In section 2, we discuss the methods of graph embedding and review the closely related work on disentangled representation learning in the context of the graph auto-encoder. Section 3 provides an introduction to the preliminaries and the problem formulation. In section 4, we propose our models, including DVGA and DGA, along with their optimization objectives and algorithms. In section 5, we derive the algorithm’s time complexity and analyze the number of algorithm’s parameter. In section 6, we present quantitative and qualitative experiments as well as model ablation, convergence behavior, and complexity analysis. The conclusion in the final section emphasizes the innovation of our approach and suggests potential directions for future research.
2 Related Work
2.1 Graph Embedding
@@
The goal of graph embedding is to embed graph-structured data into a low-dimensional feature space while preserving the graph information. We classify graph embedding algorithms into two main types: topological embedding methods and content-enhanced embedding methods (Cai et al., 2018; Pan et al., 2018).
@@
Topological embedding methods only utilize the topological structure of the graph, aiming to maximally preserve the topological information. The DeepWalk model, introduced by Perozzi et al. (2014), learned node embeddings by random walks. Later, various probabilistic models have been proposed, such as LINE (Tang et al., 2015) and node2vec (Grover et al., 2016). Matrix factorization methods like HOPE (Ou et al., 2016) and M-NMF (Wang et al., 2017) have been developed to learn the latent representation of the graph, which leveraged the adjacency matrix. In the work by Xingyi et al. (2021), node representations were generated by learning proximity measures and singular value decomposition decomposition parameters within a comprehensive framework. Deep learning models (Zaiqiao et al., 2019) have also been employed for graph embedding, such as SDNE (Wang et al., 2016) and DNGR (Cao et al., 2016).
@@
Content-enhanced embedding methods leverage both topological information and content features. Yang et al. (2015) introduced TADW which extends its matrix factorization model to support embedding of attribute information. UPP-SNE (Zhang et al., 2017) employed an approximate kernel mapping method to improve the learning of user embeddings in social networks using user profile features. Deep learning models such as GCN (Kipf et al., 2016b), GraphSAGE (Hamilton et al., 2017), GAT (Velickovic et al., 2017; Brody et al., 2021) and GSN (Bouritsas et al., 2022) can also learn node representations through supervised tasks. Some works (Kipf et al., 2016a; Simonovsky et al., 2018) were based on generative models, specifically using Variational Auto-Encoder (VAE) (Kingma et al., 2013) on graphs for learning representations. A comprehensive discussion on this topic will be provided in the subsequent subsection.
2.2 Extended (Variational) Graph Auto-Encoder
@@
With the rapid development of deep learning, there has been an increasing focus on research in graph representation learning using generative models. Inspired by AE and VAE (Kingma et al., 2013), Variational Graph Auto-Encoder (VGAE) and the Graph Auto-Encoder (GAE) (Kipf et al., 2016a) utilized GCN as encoders and inner products as decoders. They accomplished link prediction tasks while learning embeddings for each node in the graph.
@@
In recent years, significant progress has been made in enhancing the structure of these models. LGAE (Salha et al., 2021) employed significantly simpler and more interpretable linear models as encoder, achieving comparable performance across various real-world datasets. NF-VGA (Shan et al., 2020) designed a prior-generative module to generate a flexible distribution as the prior for latent representations. SIG-VAE (Hasanzadeh et al., 2019) enhanced the generative modeling by employing a hierarchical variational framework.
In terms of decoders, TVGA (Shi et al., 2020) leveraged the triadic closure property to introduce a triad decoder. The triad decoder collectively predicts the three edges within a local triad to improve link prediction tasks. Graphite (Grover et al., 2019) employed an innovative iterative refinement strategy for graphs, inspired by low-rank approximations. In terms of model training improvements, ARGA and ARVGA (Pan et al., 2018) introduced a novel adversarial graph embedding method.
@@
Recently, there have also been additional improvements from the point of refining priors and other aspects of the model (Ahn et al., 2021; Davidson et al., 2018; Sun et al., 2021; Khan et al., 2021). However, current approaches based on VGAE or GAE have only explored general settings in handling entanglement situations. Their encoders, lacking the integration of disentangled representation learning, are challenged by recognizing and disentangling the heterogeneous latent factors hidden in the observed graph-structured data. Thus, these holistic encoding approaches exhibit limited capabilities in learning graph representations and often lead to less satisfactory performance in link prediction tasks and other downstream applications.
2.3 Disentangled representation learning
Learning disentangled representation is to obtain a factorized representation which can effectively identify and disentangle the underlying explanatory factors in observed data (Bengio et al., 2013). Due to its capability to generate robust and interpretable representations, disentangled representation learning has become a central challenge in the field of machine learning.
@@ 2.3 Disentangled representation learning
The majority of existing research can be found in the field of computer vision, where variational methods based on VAE are widely employed for disentangled representation learning in images (Burgess et al., 2018; Higgins et al., 2016; Chen et al., 2018; Trauble et al., 2021; Kim et al., 2018; Locatello et al., 2020; Kim et al., 2019; Shen et al., 2022). This is primarily achieved by imposing independent constraints on the posterior of the latent variable through KL divergence.
@@
The exploration of disentangled representation learning for graph-structured data has recently captured significant attention. In many cases, the relationships between nodes in a homogeneous graph remain heterogeneous, entangled, but they are solely represented as single binary-valued edges. However, edges in the graph often carry extensive relationship information beyond binary indicators of structural connectivity. So this motivation prompts us to implicitly discover the underlying relationships between entities. DisenGCN (Ma et al., 2019) stood out as an innovative contribution in graph disentanglement. It focuses on the generation of independent latent features to achieve disentanglement at the node level. However, it did not take into account multiple potential relationships between entities. Later, ADGCN (Zheng et al., 2021) considered the existence of multiple relationships between nodes and utilized an adversarial regularizer to enhance separability among different latent factors. Related works mainly concentrate on node-level (Guo et al., 2022; Liu et al., 2020), edge-level (Zhao et al., 2022; Wu et al., 2021), and graph-level (Wu et al., 2022; Yang et al., 2015) disentanglement representation learning. Furthermore, another line of work employs a self-supervised learning strategy to obtain disentangled representations (Xiao et al., 2022; Li et al., 2021; Zhang et al., 2023). This includes the design of heuristic proxy tasks and integration with contrastive learning. Simultaneously, there has been a considerable amount of research on disentangled representation learning for heterogeneous graph (Wu et al., 2021; Wang et al., 2020; Geng et al., 2022). Nevertheless, these works either typically rely on supervised labels and primarily focus on node classification tasks, or they concentrate on other specific application domains. In contrast, our emphasis is on utilizing generative models to learn disentangled node embedding for graph-structured data, and focusing on link prediction and other downstream tasks.
@@
Currently, the exploration of disentangled representation learning for graphs through generative models, i.e. (variational) graph auto-encoder, is largely unexplored. There is only one nearly contemporaneous work that addresses this, but only focusing on link prediction task (Fu et al., 2023). In their proposed model, they incorporate an neighborhood routing mechanism (Ma et al., 2019) to achieve disentangled node embeddings. However, a drawback is that it assumes relationships between nodes are only influenced by one single factor. On the contrary, our models not only employ a more general mechanism to learn representations but also enhances the inference model’s capability through flow models. Additionally, we use a factor-wise decoder that better aligns with the link prediction task emphasized in generative models. Our objective is to learn disentangled node representations. This not only improves performance in link prediction tasks but also demonstrates outstanding results in other downstream tasks.
3 Preliminaries and Problem Formulation

### v2

@@ Abstract
The (variational) graph auto-encoder is widely used to learn representations for graph-structured data. However, the formation of real-world graphs is a complicated and heterogeneous process influenced by latent factors. Existing encoders are fundamentally holistic, neglecting the entanglement of latent factors. This reduces the effectiveness of graph analysis tasks, while also making it more difficult to explain the learned representations. As a result, learning disentangled graph representations with the (variational) graph auto-encoder poses significant challenges and remains largely unexplored in the current research. In this paper, we introduce the Disentangled Graph Auto-Encoder (DGA) and the Disentangled Variational Graph Auto-Encoder (DVGA) to learn disentangled representations. Specifically, we first design a disentangled graph convolutional network with multi-channel message-passing layers to serve as the encoder. This allows each channel to aggregate information about each latent factor. The disentangled variational graph auto-encoder’s expressive capability is then enhanced by applying a component-wise flow to each channel. In addition, we construct a factor-wise decoder that takes into account the characteristics of disentangled representations. We improve the independence of representations by imposing independence constraints on the mapping channels for distinct latent factors. Empirical experiments on both synthetic and real-world datasets demonstrate the superiority of our proposed method compared to several state-of-the-art baselines.
Index Terms:
@@
Graph auto-encoder, graph neural network, disentangled representation learning, normalizing flow.
publicationid: pubid: $0$–$00$ 2021 IEEE
I Introduction
@@
Graph-structured data has become exceptionally prevalent in the era of big data, spanning across various domains, including social networks, co-authorship networks, biological networks and traffic networks, etc. Represented in the form of graphs, these data illustrate the complex interconnections between entities, reflecting intricate and extensive networks of relationships.
The field of graph analysis includes many significant tasks. Link prediction [1, 2, 3] is a well-known task in which missing or potential connections in a network are predicted. For example, in social networks, link prediction facilitates precise targeting for target marketing, hence informing marketing strategies. Biological networks use it to forecast associations between molecules, whereas terrorist networks use it to identify individuals who arouse suspicion. Additional prevalent graph analysis tasks include node clustering [4, 5] and node classification [6, 7, 8]. Node clustering is the process of grouping nodes in a graph with the goal of ensuring that nodes inside each cluster are closely related. Node classification involves assigning predefined labels to nodes based on their attributes, and this work is essential for gaining insights into the specific characteristics of each node in a graph. These tasks also have wide-ranging applications in areas such as social networks, biology, network security and medical analysis.
However, due to the extensive and highly sparse nature of graphs, graph analysis tasks are often challenging. Therefore, there has been a growing interest in improving the learning of graph representations [9, 10, 11]. The tasks of learning node representations serve as the basis for the previously mentioned network analysis tasks. An effective approach for learning node representations without supervision is the Variational Graph Auto-Encoder (VGAE) and the Graph Auto-Encoder (GAE) [12]. They all consist of an encoder responsible for learning latent representations and a decoder tasked with reconstructing the graph structure data using the learned representations. Their primary objective is to successfully complete the task of link prediction. Consequently, the decoder is relatively simple, depending only on the inner product of the latent representations of the nodes. The encoder is based solely on the original Graph Convolutional Networks (GCN) [13]. These methods take a holistic approach and treat the graph as a perceptual whole [12, 14, 15, 16, 17, 18, 19]. They ignore subtle differences between various relationships in the graph. In fact, graph formation in the real world is a complex, heterogeneous process driven by latent factors. Figure 1 illustrates the concept of disentanglement. It is evident that the connections between the four persons are established based on several underlying factors. Moreover, their relationships exhibit distinct emphases. For instance, the relationship between person $A$ and person $B$ is mainly due to familial connections, whereas the connection between person $A$ and person $C$ is predominantly based on their shared hobby. Additionally, it is important to acknowledge that relationships between any pair of individuals are influenced by multiple factors, while the emphasis placed on these aspects may differ. Without this precise differentiation in relationships, representing nodes only from binary network topologies becomes problematic. This not only hinders the interpretability of learned representations but also results in diminished performance on subsequent tasks. Hence, the intricate nature of the problem forces the encoder to disentangle factors, a crucial aspect that current holistic methods have overlooked.
@@
In this paper, we propose learning disentangled node representations. Disentangled representation learning, which seeks to identify different underlying factors in factorized representations, has shown many advantages [20]. However, it encounters three problems when applied to graph auto-encoder models. (1) Designing a graph encoder for disentangled representations requires careful planning to guarantee it can effectively infer latent factors with enough expressiveness. (2) Designing a decoder for disentangled representations requires taking into account the factor information in the connections between nodes in order to enhance the accuracy of link prediction. (3) To improve the quality of disentangled representations, a training strategy should be designed to ensure statistical independence between distinct parts of representations that correspond to different latent factors.
Figure 1: An illustrative social network example that inspired our work. We consider using lines of different colors to represent different underlying factors that contribute to the relationship between two persons. The thickness of the lines is directly proportional to the frequency of interaction due to these factors.
@@
To tackle the aforementioned challenges, we introduce two innovative auto-encoder models for learning disentangled graph representations, that is, Disentangled Variational Graph Auto-Encoder (DVGA) and Disentangled Graph Auto-Encoder (DGA). Initially, we design a Disentangled Graph Convolutional Network (DGCN) employing dynamic assignment mechanism [21], incorporating multiple channels to aggregate information related to each disentangled latent factor. Then, a component-wise flow is applied to each channel to enhance the expressiveness. Furthermore, we design a factor-wise decoder, ensuring that if any factor contributes to a link between nodes, then those two nodes will be connected in the final prediction result. To encourage the independence of representations corresponding to different latent factors, we impose independence constraints on the mapping channels of different latent factors. Finally, the auto-encoder and independent regularization will be co-optimized in a unified framework, allowing the disentangled graph encoder to learn better disentangled graph representations and improve the overall performance of the model. Our models differ from existing methods by learning disentangled representations from the graph. This enables us to explore the significance of each channel, leading to enhanced interpretability and the generation of persuasive and comprehensible link prediction results.
@@
  1. 1.
    We present two novel disentangled graph auto-encoder models: the Disentangled Variational Graph Auto-Encoder (DVGA) and the Disentangled Graph Auto-Encoder (DGA). The DVGA model incorporates a disentangled encoder, component-wise flow, factor-wise decoder, and an independence regularization term in the loss function to learn disentangled representations. Similarly, the DGA model is also introduced. As far as we know, we are one of the only researchers studying disentangled graph auto-encoder models.
  2. 2.
    Our proposed disentangled graph convolutional network in the encoder captures multiple factors by learning disentangled latent factors on graphs, enabling multi-relation disentangling. We enhance the expressive capability of each channel representation by incorporating flow, hence enhancing the overall capability of the representation. The designed decoder is tailored for disentangled representations, improving its ability to predict graph structures.
  3. 3.
@@
II Related Work
In this section, we give a brief review of the research topics that are relevant to our proposed method, including graph embedding, extended (variational) graph auto-encoder, and disentangled representation learning.
II-A Graph Embedding
@@
The goal of graph embedding is to embed graph-structured data into a low-dimensional feature space while preserving the graph information. We classify graph embedding algorithms into two main types: topological embedding methods and content-enhanced embedding methods [22, 23].
@@
Topological embedding methods only utilize the topological structure of the graph, aiming to maximally preserve the topological information. The DeepWalk model, introduced by [24], learned node embeddings by random walks. Later, various probabilistic models have been proposed, such as LINE [25] and node2vec [26]. Matrix factorization methods like HOPE [27] and M-NMF [10] have been developed to learn the latent representation of the graph, which leveraged the adjacency matrix. Deep learning models [28] have also been employed for graph embedding, such as SDNE [9] and DNGR [29].
@@
Content-enhanced embedding methods leverage both topological information and content features. TADW [30] extended the matrix factorization model to support embedding of attribute information. UPP-SNE [31] employed an approximate kernel mapping method to improve the learning of user embeddings in social networks using user profile features. Deep learning models such as GCN [13], GraphSAGE [32], GAT [33] and GSN [34] learned node representations through supervised tasks. Some works [12, 35] were based on generative models, specifically using Variational Auto-Encoder (VAE) [36] on graphs for learning representations. A comprehensive discussion on this topic will be provided in the subsequent subsection.
II-B Extended (Variational) Graph Auto-Encoder
@@
With the rapid development of deep learning, there has been an increasing focus on research in graph representation learning using generative models. Inspired by AE and VAE [36], Variational Graph Auto-Encoder (VGAE) and the Graph Auto-Encoder (GAE) [12] utilized GCN as encoders and inner products as decoders. They accomplished link prediction tasks while learning embeddings for each node in the graph.
@@
In recent years, significant progress has been made in enhancing the structure of these two models. LGAE [37] employed significantly simpler and more interpretable linear models as encoder. NF-VGA [38] designed a prior-generative module to generate a flexible distribution as the prior for latent representations. SIG-VAE [14] enhanced the generative modeling by employing a hierarchical variational framework. In terms of decoder, TVGA [17] leveraged the triadic closure property to introduce a triad decoder. Graphite [15] employed an innovative iterative refinement strategy for graphs, inspired by low-rank approximations. In terms of model training improvements, ARGA and ARVGA [23] introduced a novel adversarial graph embedding method.
@@
There have also been other breakthroughs in terms of refining priors and other components of the models [39, 18, 16]. However, current approaches based on VGAE or GAE have only explored general settings in handling entanglement situations. Their encoders, lacking the integration of disentangled representation learning module, are challenged in recognizing and disentangling the heterogeneous latent factors hidden in the observed graph-structured data. Thus, these holistic encoding approaches exhibit limited capability in learning graph representations and often lead to less satisfying performance in link prediction tasks and other downstream applications.
II-C Disentangled representation learning
@@
The goal of learning disentangled representation is to obtain a factorized representation that can effectively identify and disentangle the underlying factors in observed data [40]. Disentangled representation learning has become a prominent challenge in machine learning due to its ability to generate robust and interpretable representations. The majority of existing research can be found in the field of computer vision, where variational methods based on VAE are widely employed for disentangled representation learning in images [41, 42, 43, 44]. This is mainly achieved by imposing independent constraints on the posterior distribution of the latent variable through KL divergence.
@@
There has been a recent surge of interest in studying disentangled representation learning for graph-structured data. The connections between nodes in a homogeneous graph remain heterogeneous and entangled, but they are only represented as single binary-valued edges. However, the edges in the graph often contain substantial relationship information that goes beyond simple binary indicators of structural connectivity. This motivation prompts us to implicitly uncover the underlying relationships between entities. DisenGCN [45] is a notable and original advancement in the field of graph disentanglement. However, it did not take into account multiple potential relationships between entities. Later, ADGCN [21] considered the existence of multiple relationships between nodes and utilized an adversarial regularizer to enhance separability among different latent factors. Related works mainly focused on node-level [46, 47], edge-level [48, 49], and graph-level [50, 51] disentanglement representation learning. Furthermore, another line of work employed a self-supervised learning strategy to learn disentangled representations [52, 53, 54]. Simultaneously, there has been a significant amount of study on disentangled representation learning for heterogeneous graphs [55, 56]. Nevertheless, these works either typically rely on supervised labels and largely focus on node classification tasks, or they concentrate on other specific application domains. In contrast, we emphasize the use of generative models to learn disentangled node embedding for graph-structured data and focus on link prediction and other downstream tasks.
@@
Currently, there is limited research on the study of disentangled representation learning for graphs using generative models, specifically (variational) graph auto-encoder. There is just one study that is almost contemporary but specifically deals with the task of link prediction, as mentioned in [57]. The authors employed a neighborhood routing method [45] in their suggested model to obtain disentangled node embeddings. However, a drawback of this approach is its assumption that interactions between nodes are solely influenced by a single factor. In contrast, our models not only utilize a more general mechanism for learning representations but also boost the inference model’s capacity through flow models. In addition, we construct a factor-wise decoder that is more closely aligned with the link prediction task that is stressed in generative models. The aim of our study is to learn disentangled node representations. Not only does this enhance performance in link prediction, but it also demonstrates excellent results in other subsequent downstream tasks.
III Preliminaries and Problem Formulation